OpenStudy (anonymous):

Which of the following best describes the graph of y=−x2? A.The domain only includes numbers greater than or equal to 0. B.The range only includes numbers greater than or equal to 0. C.Both the domain and range only include numbers greater than or equal to 0. D.The domain and range include all real numbers

OpenStudy (anonymous):

That's the graph of y = x^2, what do you think y = -x^2 will look like? What will happen if I stick a negative sign in front of that?

OpenStudy (anonymous):

And btw, neither A, B, C, or D is correct.